# The importance of visual modelling languages in generative software engineering

### v1

@@ 4.4 Hand drawn diagrams
Copilot correctly recognises the fact that that the activity diagram represents a sorting algorithm, and returns a method implementing it.
@@ 4.5 Design patterns
Design Patterns Gamma et al. (1994) are reusable solutions to commonly occurring problems in software design. They are not specific implementations but rather general templates that can be adapted to various situations. By using design patterns, developers can create more flexible, maintainable, and efficient software systems. There exists a wealth of existing patterns available in the SE literature; In addition, new patterns can be created and illustrate via appropriate UML diagrams. For instance, we may consider the “Adapter” pattern, which converts the interface of a class into another interface clients expect. This pattern is illustrated in Figure 12.
@@
In this system we have two portable drives, a Samsung drive and a Western Digital drive. The Samsung drive features a microusb connector, while the Western Digital drive features a traditional usb connector. Our aim is to allow the Samsung drive to read and write data via usb. To achieve this, we attach the diagram to Copilot and provide the following prompt.
@@
5 Discussion
Throughout Section 4 we have utilised Microsoft Copilot in its web-based version, which is based on GPT-4o and allows image attachments. To ensure reproducibility, we have also developed an equivalent pipeline in Gemini, formalised in a Jupyter Notebook based on gemini-1.5-flash, which is included in our SM.
@@ 5 Discussion
In this section, we reflect on how the use cases discussed in Section 4, which are made possible thanks to the enhanced capabilities of GPT-4, address several research gaps in GenSE, as well as some long standing issues in SE.
@@
We shall focus first on the research gaps we bridged in GenSE. First, our study addresses the lack of applications of GenAI to the software design sub-task of SE, a literature gap identified in Huang et al. (2024); Hou et al. (2024). Second, to the best of our knowledge, our study represents the first application of multimodal GPTs (diagram + prompt) to the software development sub-task of SE, which in GenSE is generally carried out via natural language-based prompts. Third, we have discussed and illustrated with an empirical example the superiority, in terms of information transfer efficiency, of diagram-based prompting against pure natural language prompting; a result that aligns with the discussion in Tufte (2003). Fourth, by reverse engineering code into diagrams, we have contributed to the code summarisation sub-task of SE, which is again predominantly carried out via natural language summaries. Fifth, while there exist studies that investigated the production of PlantUML code from hand drawn diagrams Conrardy and Cabot (2024); Wang et al. (2024), we believe this is the first study in which we cover multiple elements of the SE stack, from conceptual modelling — which includes static and dynamic analysis of the software system — to implementation. Finally, in addition, the use of existing (or new) design patterns in the context of Few-Shot prompting for software system design does not seem to have been investigated before in the literature.
@@
Next, we shall look at which long standing issues in SE may be addressed by the use cases discussed in our work. In Baltes and Diehl (2014), the authors speculated that, as documentation is frequently poorly written and out of date, sketches may serve as a supplement to conventional documentation like source code comments. Consequently, they advocated the development of suitable tools to support archiving and retrieval of such sketches. No such tool has been developed ever since and poorly documented code remains a persistent issue in SE. However, GenAI has already started to change this picture via automatic generation of code comments. And yet, if sketches and diagram (both formal and informal) can be transformed, as we have shown, into a preliminary draft of the desired source code, then producing reliable visual descriptions of the software system under development suddenly becomes appealing. Likewise, if it is possible to selectively and cheaply reverse engineer part of such software system to visually inspect the behaviour of classes and methods, the developer becomes equipped with a formidable arsenal of high resolution communication tools to enhance code understanding and boost team communication.

### v2

@@ 4.4 Hand-drawn diagrams
Copilot correctly recognises the fact that that the activity diagram represents a sorting algorithm (bubble sort), and returns a method implementing it.
@@ 4.5 Design patterns
Design Patterns Gamma et al. (1994) are reusable solutions to commonly occurring problems in software design. They are not specific implementations but rather general templates that can be adapted to various situations. By using design patterns, developers can create more flexible, maintainable, and efficient software systems. There exists a wealth of existing patterns available in the SE literature; In addition, new patterns can be created and illustrated via appropriate UML diagrams. For instance, we may consider the “Adapter” pattern, which converts the interface of a class into another interface clients expect. This pattern is illustrated in Figure 12.
@@
Consider the class diagram in Figure 13. In this system we have two portable drives, a Samsung drive and a Western Digital drive. The Samsung drive features a microusb connector, while the Western Digital drive features a traditional usb connector. Our aim is to allow the Samsung drive to read and write data via usb. To achieve this, we attach the diagram to Copilot and provide the following prompt.
@@
5 Discussion and conclusions
@@
In this section, we reflect on how the use cases discussed in Section 4, which are made possible thanks to the enhanced capabilities of GPT-4, address several research gaps in GenSE, as well as some long standing issues in SE. We will also discuss how the same use cases open new research directions that can be investigated in future studies.
@@
We shall focus first on the research gaps we bridged in GenSE. First, our study addresses the lack of applications of GenAI to the software design sub-task of SE, a literature gap identified in Huang et al. (2024); Hou et al. (2024). Second, to the best of our knowledge, our study represents the first application of multimodal GPTs (diagram + prompt) to the software development sub-task of SE, which in GenSE is generally carried out via natural language-based prompts. Third, we have discussed and illustrated with an empirical example that diagram-based prompting may be advantageous, in terms of information transfer efficiency, compared to pure natural language prompting; a result that aligns with the discussion in Tufte (2003). Fourth, by reverse engineering code into diagrams, we have contributed to the code summarisation sub-task of SE, which is again predominantly carried out via natural language summaries. Fifth, while there exist studies that investigated the production of PlantUML code from hand-drawn diagrams Conrardy and Cabot (2024); Wang et al. (2024), we believe this is the first study in which we cover multiple elements of the SE stack, from conceptual modelling — which includes static and dynamic analysis of the software system — to implementation. Finally, in addition, the use of existing (or new) design patterns in the context of Few-Shot prompting for software system design does not seem to have been investigated before in the literature, and hence merits attention in future studies.
@@
Next, we shall look at which long standing issues in SE may be addressed by the use cases discussed in our work. In Baltes and Diehl (2014), the authors speculated that, as documentation is frequently poorly written and out of date, sketches may serve as a supplement to conventional documentation like source code comments. Consequently, they advocated the development of suitable tools to support archiving and retrieval of such sketches. No such tool has been developed ever since and poorly documented code remains a persistent issue in SE. However, GenAI has already started to change this picture via automatic generation of code comments. And yet, if sketches and diagram (both formal and informal) can be transformed, as we have shown, into a preliminary draft of the desired source code, then producing reliable visual descriptions of the software system under development suddenly becomes appealing. Likewise, if it is possible to selectively and cheaply reverse engineer part of such software system to visually inspect the behaviour of classes and methods, the developer becomes equipped with a formidable arsenal of high resolution communication tools to enhance code understanding, boost team communication, and resolve the long lasting conflict between investing time in developing working software or writing comprehensive documentation. These are all important and novel directions of enquiry that should be investigated in future studies.